OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 2. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

IS it Log base 16 of 32?

OpenStudy (anonymous):

yes. log base 16 of 32 = x + 2

OpenStudy (anonymous):

If it is : \[\log_{16} 32 = \log_{16}(16\times2)=\log_{16}16 + \log_{16}2 = 1 + [\log_{2}2/\log_{2}16] = 1 + 1 / [4\log_{2}2] \]

OpenStudy (anonymous):

\[1 + 1/[4\log_{2}2] = 1 + 1/4 = 5/4 \]

OpenStudy (anonymous):

So know you have 5 / 4 = x+ 2 5 = 4x + 8 4x = -3 x = -3/4

OpenStudy (anonymous):

The main part of this question is using property of logs to simplify the left side of the equation to 5/4